i) Let S={2,4,7} and T={1,3,5}. Find f(S×T) if

f(x,y)=⌊14x/3y⌋


Type/Insert your answer here!

Note: No partial credit would be admissible in this question


f(x,y)=x^2+y^3


Type/Insert your answer here!

Note: No partial credit would be admissible in this question


1
Expert's answer
2021-05-07T09:25:46-0400

S={2,4,7}, T={1,3,5}


S"\times" T={(2,1),(2,3),(2,5),(4,1),(4,3),(4,5),(7,1),(7,3),(7,5)}


(i) "f(x,y)=\dfrac{14x}{3y}"


So,

"f(S\times T)="

"\dfrac{14(2)}{3(1)}+\dfrac{14(2)}{3(3)}+\dfrac{14(2)}{3(5)}+\dfrac{14(4)}{3(1)}+\dfrac{14(4)}{3(3)}+\dfrac{14(4)}{3(5)}+\dfrac{14(7)}{3(1)}+\dfrac{14(7)}{3(3)}+\dfrac{14(7)}{3(5)}"


Hence "f(S\times T)" ={"\dfrac{28}{3},\dfrac{28}{9},\dfrac{28}{15},\dfrac{56}{3},\dfrac{56}{9},\dfrac{56}{15},\dfrac{98}{3},\dfrac{98}{9},\dfrac{98}{15}" }


(ii) "f(x,y)=x^2+y^2"


"f(S\times T)" ={

"(2^2+1^2),(2^2+3^2),(2^2+5^2),(4^2+1^2),(4^2+3^2),(4^2+5^2),\\(7^2+1^2),(7^2+3^2),(7^2+5^2)" }


"F(S\times T)=" {"{5,13,29,17,25,41,50,58,74}" }